# VertiBench: Advancing Feature Distribution Diversity in Vertical Federated Learning Benchmarks

### v1

@@ Appendix C Split Method Details
       $k\leftarrow\text{Random choice from }[1,K]\cup\mathbb{N}\;\;w.p.\;\;(p_{1},\dots,p_{K})$;
@@ D.1 Correlation Evaluation Metric
Our observations, as presented in Figure 3(a) and 3(b), reveal the following: (1) Pcor behaves similarly to mcor when evaluating inner-party correlation and shows a similar trend to mcor Taylor [2020] when assessing inter-party correlation. (2) Both Pcor and mcor exhibit the lowest inter-party correlation and the highest inner-party correlation at the extremities of the x-axis. This suggests that the datasets $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are managed by distinct parties and exhibit complete independence. This pattern is also reflected in Figure 3(c) when Pcor is applied to parties of different dimensions. These observations validate the appropriateness of Pcor as a measure for evaluating inter-party correlation, even when the number of features between the two parties varies.
@@
D.2 Correlation-based Feature-split Algorithm
We turn our attention towards validating the efficacy of our proposed correlation-based feature-split algorithm. Three synthetic datasets, each encompassing 10 features, are independently generated using the sklearn library. These datasets are subsequently concatenated along the feature axis, yielding a global dataset with 30 features. This dataset is then split into three local datasets, each containing 10 features, deploying our proposed algorithm with $\beta$ values set at 0, 0.5, and 1.0.
The visualization of the absolute value of the correlation matrix between each pair of features is presented in Figure 5. As evident in Figure 4(a), 4(b), and 4(c), an increment in the value of $\beta$ corresponds to an increase in the correlation between features of different parties. This observation is in alignment with our expectations. However, randomly splitting features into three parties, serving as our baseline, fails to reflect scenarios with low inter-party correlation (as illustrated in Figure 4(d)). These empirical findings underscore the efficacy of our proposed algorithm in partitioning features based on their correlation attributes.
@@ D.2 Correlation-based Feature-split Algorithm
(d) random-split
Figure 5: Correlation matrix of the global dataset with party boundaries indicated by red lines
D.3 Time Efficiency of Split Methods.
@@ Appendix E Real Dataset Construction
In this section, we outline the construction process of the Satellite dataset, which was adapted from the WorldStrat dataset Cornebise et al. [2022], originally intended for high-resolution imagery analysis. The Satellite dataset encompasses Point of Interest (POI) data, each associated with one or more Areas of Interest (AOI). Every AOI incorporates a unique location identifier, a land type, and 16 low-resolution, 13-channel images, each taken during a satellite visit to the location.
@@
The Satellite dataset forms a practical VFL scenario for location identification based on satellite imagery. Each AOI, with its unique location identifier, is captured by 16 satellite visits. Assuming each visit is carried out by a distinct satellite organization, these organizations aim to collectively train a model to classify the land type of the location without sharing original images. The Satellite dataset encompasses four land types as labels, namely Amnesty POI (4.8%), ASMSpotter (8.9%), Landcover (61.3%), and UNHCR (25.0%), making the task a 4-class classification problem of 3927 locations.
@@
Our use of the WorldStrat dataset was restricted to the labels and Sentinel2 imagery, falling under the CC BY 4.0 lic [2023a] license, while excluding high-resolution imagery that falls under the CC BY-NC 4.0 lic [2023b] license. Therefore, we have released the Satellite dataset under the CC BY 4.0 license.
@@
We aim to create a dedicated website for federated learning datasets to host the Satellite dataset and future VFL datasets. Although the website is currently under construction, we have made the Satellite dataset available via a public Google Drive link Anonymized [2023] for review purposes. The provided ZIP file comprises 32 CSV files, corresponding to training and testing datasets split at a ratio of 8:2. Each training and testing file contains 3,142 and 785 flattened images from a party, respectively.
@@ Appendix F Experimental Details
The datasets employed in our experiments exhibit a range of dimensions (from 16 to 20,958), instance numbers (from 15k to 581k), and tasks, which include binary classification, multi-class classification, and regression. Detailed information about these datasets and the corresponding licenses are presented in Table 4.
Table 4: Detailed information of datasets (N/A means unspecified)
Dataset Task #samples #features #classes License
covtype dat [1998] classification 581,012 54 7 BSD lic [2023c]
msd dat [2011] regression 463,715 90 1 CC BY-NC-SA 2.0 lic [2023d]
gisette dat [2008a] classification 60,000 5,000 2 CC BY 4.0 lic [2023a]
realsim dat [2015] classification 72,309 20,958 2 N/A
epsilon dat [2008b] classification 400,000 2,000 2 N/A
letter dat [1991] classification 15,000 16 26 N/A
radar Khosravi [2020] classification 325,834 174 7 N/A
@@
With regard to Split-NN-based models, specifically SplitNN and C-VFL, each local model is trained by a two-layer multi-layer perceptron (MLP) with each hidden layer containing 100 units. The corresponding aggregated model is a single-layer MLP with 200 hidden units. The learning rate, chosen from the set $\{10^{-4},10^{-3},3\times 10^{-3}\}$, is contingent on the specific algorithm and dataset. The number of iterations is fixed at 50 for SplitNN and 200 for C-VFL, with the latter setting aimed at ensuring model convergence. We also test C-VFL using four quantization buckets, a single vector quantization dimension, and a top-k compressor as recommended in the default setting. The number of local rounds $Q$ in C-VFL is set to 10.
@@
The hardware configuration used for C-VFL, GAL, SplitNN, and FedTree consists of 2x AMD EPYC 7543 32-Core Processors, 4x A100 GPUs, and 503.4 GB of RAM, running on Python 3.10.11 with PyTorch 2.0.0, Linux 5.15.0-71-generic, Ubuntu 22.04.2 LTS. For FATE framework, we are using federatedai/standalone_fate Docker image, running with Python 3.8.13 on Docker 23.0.2. Pivot is compiled from source using CMake 3.19.7, g++ 9.5.0, libboost 1.71.0, libscapi with git commit hash 1f70a88, and runs on a slurm cluster with AMD EPYC 7V13 64-Core Processor with the same number of cores as 2x AMD EPYC 7543 used for other algorithms.
@@
GAL Diao et al. [2022a] MIT lic [2023e]
C-VFL Castiglia et al. [2022a] MIT lic [2023e]
SecureBoost Cheng et al. [2021a] Apache 2.0 lic [2023f]
Pivot Wu et al. [2020a] CC BY-NC-ND 4.0 lic [2023g]
FedTree Li et al. [2023] Apache 2.0 lic [2023f]
  • Note: AL Xian et al. [2020a], SplitNN Vepakomma et al. [2018], BlindFL Fu et al. [2022b] and VF2Boost Fu et al. [2021] do not specify their licensing terms.
@@ G.1 Communication Cost Details
In this subsection, we analyze the maximum incoming and outgoing communication of each VFL algorithm (Figure 6), which aligns with our observations in Section 6.3.
@@
Generally, GAL is more efficient than FedTree, albeit FedTree has less incoming communication cost on low-dimensional datasets. The incoming communication cost includes residuals for GAL and histograms for FedTree. Residual size is dependent on the number of instances, while histogram size relates to the feature count. Hence, FedTree shows less incoming communication cost on low-dimensional datasets but more on high-dimensional datasets. However, FedTree’s outgoing communication cost is approximately twice that of GAL, as it transmits both gradients and Hessians, neutralizing its incoming communication efficiency on low-dimensional datasets.
C-VFL primarily reduces communication costs via representation compression sent to the server, yet it does not enhance the backward server-client communication. This substantial reduction in the incoming communication cost is indicative of a noteworthy decrease in the representation size. Conversely, the consistent outgoing communication cost suggests that the transmission of compressed aggregated models within C-VFL is not demonstrating greater efficiency compared to the process of transmitting uncompressed gradients of the cut layer. Such an observation provides insights into potential enhancements that could be pursued to reduce SplitNN’s backward communication costs.
@@
Figure 6: The max incoming and outgoing communication size of VFL algorithms (50 global iterations)
@@ G.3 Training Time
Secondly, when comparing non-encryption methods, we find that split-based algorithms (SplitNN, FedTree) generally outperform ensemble-based algorithms (GAL) in terms of efficiency. This is primarily because split-based algorithms require each party to train a partial model, whereas ensemble-based algorithms mandate that each party train an entire model for ensemble purposes. This design characteristic also contributes to the lower communication costs associated with ensemble-based algorithms, as demonstrated in Figure 3 and Figure 6.

### v2

@@ Appendix C Split Method Details
       $k\leftarrow\text{Random choice from }[1,K]\cap\mathbb{N}\;\;w.p.\;\;(p_{1},\dots,p_{K})$;
@@ D.1 Correlation Evaluation Metric
Our observations, as presented in Figure 4(a) and 4(b), reveal the following: (1) Pcor behaves similarly to mcor when evaluating inner-party correlation and shows a similar trend to mcor Taylor [2020] when assessing inter-party correlation. (2) Both Pcor and mcor exhibit the lowest inter-party correlation and the highest inner-party correlation at the extremities of the x-axis. This suggests that the datasets $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are managed by distinct parties and exhibit complete independence. This pattern is also reflected in Figure 4(c) when Pcor is applied to parties of different dimensions. These observations validate the appropriateness of Pcor as a measure for evaluating inter-party correlation, even when the number of features between the two parties varies.
@@
D.2 Feature-split Algorithm
Importance-based Feature Split.
We rigorously investigate the correlation between $\{\alpha_{i}\}$ and party importance, using both Shapley value and Shapley-CMI for feature importance assessment. For property (1), in a two-party VFL experiment, we fix $\alpha_{2}=1$ and vary $\alpha_{1}$ from $10^{-2}$ to $10^{3}$. For each $\alpha_{2}$, we compute the scaled $\alpha_{i}$ as $\alpha_{i}/\sum_{i}{\alpha_{i}}$ and run 1,000 trials to estimate the party importance’s expected value. The subsequent relationship between scaled $\alpha_{i}$ and the expected party importance is illustrated in Figure 5(a) and 5(b), demonstrating clear proportionality.
For property (2), four parties are divided under a symmetric Dirichlet distribution (where $\forall i,\alpha_{i}=\alpha$) with $\alpha$ values ranging from $10^{-2}$ to $10^{3}$. For each $\alpha$, the standard variance of party importance is estimated over 1,000 trials, as visualized in Figure 5(c) and 5(d). These plots signify a negative relationship between $\alpha=\|\{\alpha_{i}\}_{i=1}^{K}\|_{2}$ and the standard variance of party importance, extending the scenario of a random split leading to consistent party imbalance levels.
(a) $\alpha_{i}$ vs. Shapley
(b) $\alpha_{i}$ vs. Shapley-CMI
(c) $\alpha$ vs. Shapley std.
(d) $\alpha$ vs. Shapley-CMI std.
Figure 6: Relationship between $\alpha$ and party importance: (a)(b) Proportionality of scaled $\alpha_{i}$ to party $P_{i}$’s importance; (c)(d) Negative correlation between $\alpha$ and standard variance (std.) of party importance. The red dotted line indicates the importance variance of random split.
Correlation-based Feature Split.
We turn our attention towards validating the efficacy of our proposed correlation-based feature-split algorithm. Three synthetic datasets, each encompassing 10 features, are independently generated using the sklearn library. These datasets are subsequently concatenated along the feature axis, yielding a global dataset with 30 features. This dataset, with features shuffled, is then split into three local datasets, each containing 10 features, deploying our proposed algorithm with $\beta$ values set at 0, 0.5, and 1.0.
Specifically, we split synthetic datasets with various $\beta$ values using Algorithm 1 (as shown in Figure 6(a), 6(b), and 6(c)) and compare this to a random split (Figure 6(d)). Our findings reveal that as $\beta$ increases, inter-party correlation correspondingly escalates. Additionally, as depicted in Figure 8, when we perform splits on a concatenated feature-shuffled real VFL dataset Vehicle ($\beta=0$), VertiBench, when parameterized with the same $\beta$, exactly reconstruct the real feature split of Vehicle dataset. This indicates that VertiBench scope aligns well with the real scope which corresponds to a scenario with small inter-party correlation as further depicted in Figure 9. In contrast, a uniformly random split (uniform scope) produces a distinct correlation matrix.
@@ D.2 Feature-split Algorithm
(d) uniform-split
Figure 7: Absolute correlation matrix of the global dataset with party boundaries indicated by red lines. Icor means inter-party correlation. (a),(b),(c) - correlation-based split; (d) uniform split.
(a) Real: Vehicle dataset
(b) Synthetic: spilt with same $\beta$
(c) Synthetic: uniform split
Figure 8: Absolute correlation matrix of real Vehicle dataset and synthetic datasets generated from feature-shuffled Vehicle - VertiBench with the same $\beta$ vs. uniform split
(a) NUS-WIDE correlation
(b) Vehicle correlation
(c) Satellite correlation
Figure 9: Absolute correlation matrix of real VFL datasets with party boundaries marked in red lines
D.3 Time Efficiency of Split Methods
@@
To further evaluate the performance, we specifically arranged experiments to evaluate the computational intensity of the correlation-based splitting method. We conducted tests on synthetic datasets of varying dimensions or sizes generated by sklearn. Using a Pearson-based splitting method on a single RTX 3090 GPU, the results are detailed in Figure 10.
Scalability experiments indicate the efficiency of the proposed split method. For the importance-based split method, it avoids the computational demands of Shapley value calculations while retaining related properties. This method completed within 30 seconds across all tested datasets, ensuring scalability for large datasets. The correlation-based method effectively handles datasets with up to 10k features, aligning with the dimensional needs of common VFL datasets such as realsim and CIFAR10. Meanwhile, the number of parties hardly affect the efficiency under a fixed global dataset.
(a) Sample size impact
(#features=100)
(b) Feature size impact
(#samples=1000)
(c) Party size impact
(#samples=#parties=1000)
Figure 10: Scalability of correlation-based split algorithm on #samples and #features
@@ Appendix E Real Dataset Construction
In this section, we outline the construction process of Satellite, which was adapted from the WorldStrat dataset Cornebise et al. [2022], originally intended for high-resolution imagery analysis. The Satellite encompasses Point of Interest (POI) data, each associated with one or more Areas of Interest (AOI). Every AOI incorporates a unique location identifier, a land type, and 16 low-resolution, 13-channel images, each taken during a satellite visit to the location.
@@
Satellite forms a practical VFL scenario for location identification based on satellite imagery. Each AOI, with its unique location identifier, is captured by 16 satellite visits. Assuming each visit is carried out by a distinct satellite organization, these organizations aim to collectively train a model to classify the land type of the location without sharing original images. Satellite encompasses four land types as labels, namely Amnesty POI (4.8%), ASMSpotter (8.9%), Landcover (61.3%), and UNHCR (25.0%), making the task a 4-class classification problem of 3,927 locations.
As depicted in Figure 11, the consistency in capturing the same location 16 times presents variations in image quality. These discrepancies arise due to the changing weather and lighting conditions experienced during each satellite visit. Hence, each of these 16 satellite visits can equivalently be considered as 16 separate parties. All 13-channel images in the satellite dataset that correspond to the description of Sentinel-2 Bands as shown in Table 3.
Figure 11: Preview of an area of Satellite dataset at channel-1 with land type Amnesty
Table 3: Channel information of satellite dataset
Channel Band Description
1 Band 1 Aerosol (443 nm)
2 Band 2 Blue (490 nm)
3 Band 3 Green (560 nm)
4 Band 4 Red (665 nm)
5 Band 5 Vegetation Red Edge (705 nm)
6 Band 6 Vegetation Red Edge (740 nm)
7 Band 7 Vegetation Red Edge (783 nm)
8 Band 8 NIR (842 nm)
9 Band 8A Narrow NIR (865 nm)
10 Band 9 Water Vapor (940 nm)
11 Band 10 SWIR - Cirrus (1375 nm)
12 Band 11 SWIR (1610 nm)
13 Band 12 SWIR (2190 nm)
@@
Our use of the WorldStrat dataset was restricted to the labels and Sentinel-2 imagery, falling under the CC BY 4.0 Commons [2023a] license, while excluding high-resolution imagery that falls under the CC BY-NC 4.0 Commons [2023b] license. Therefore, we have released the Satellite dataset under the CC BY 4.0 license.
@@
We aim to create a dedicated website for federated learning datasets to host Satellite and future VFL datasets. Although the website is currently under construction, we have made Satellite available via a public Google Drive link Anonymized [2023b] for review purposes. The provided ZIP file comprises 32 CSV files, corresponding to training and testing datasets split at a ratio of 8:2. Each training and testing file contains 3,142 and 785 flattened images from a party, respectively. The code of VertiBench  Anonymized [2023a] also contains a Satellite loader for demonstration.
@@ Appendix F Experimental Details
The datasets employed in our experiments exhibit a range of dimensions (from 16 to 20,958), instance numbers (from 15k to 581k), and tasks, which include binary classification, multi-class classification, and regression. Detailed information about these datasets and the corresponding licenses are presented in Table 4 and Table 5.
Table 4: Detailed information of centralized datasets (N/A means unspecified)
[b] Dataset Task #samples #features #classes License covtype Blackard [1998] cls 581,012 54 7 BSD1 msd Bertin-Mahieux [2011] reg 463,715 90 1 CC BY-NC-SA 2.02 gisette Guyon et al. [2008] cls 6,000 5,000 2 CC BY 4.03 realsim Andrew [2015] cls 72,309 20,958 2 N/A epsilon Guo-Xun et al. [2008] cls 400,000 2,000 2 N/A letter Slate [1991] cls 15,000 16 26 N/A radar Khosravi [2020] cls 325,834 174 7 N/A MNIST Deng [2012] cls 70,000 784 10 CC BY-SA 4.0 DEED4 CIFAR10 Krizhevsky and Hinton [2009] cls 60,000 1,024 10 N/A
Table 5: Detailed information of real VFL datasets
[b] Dataset Task #parties #samples1 #features2 #classes NUS-WIDE Chua et al. [2009] cls 5 107,859 / 161,789 64 / 144 / 73 / 128 / 225 2 Vehicle Duarte and Hu [2004] cls 2 63,058 / 15,765 50 / 50 3
  • 1 The #samples is written in the form of train/test.
  • 2 The #features for each parties are divided by /.
  • The licenses of both datasets are unspecified.
@@
With regard to split-NN-based models, specifically SplitNN and C-VFL, each local model is trained by a two-layer multi-layer perceptron (MLP) with each hidden layer containing 100 units. The corresponding aggregated model is a single-layer MLP with 200 hidden units. The learning rate, chosen from the set $\{10^{-4},10^{-3},3\times 10^{-3}\}$, is contingent on the specific algorithm and dataset. The number of iterations is fixed at 50 for SplitNN and 200 for C-VFL, with the latter setting aimed at ensuring model convergence. We also test C-VFL using four quantization buckets, a single vector quantization dimension, and a top-k compressor as recommended in the default setting. The number of local rounds $Q$ in C-VFL is set to 10. In the evaluation of communication cost, the split parameter $\alpha$ is set to 0.1 since feature split hardly affects the communication size.
@@
The hardware configuration used for C-VFL, GAL, SplitNN, and FedTree consists of 2x AMD EPYC 7543 32-Core Processors, 4x A100 GPUs, and 503.4 GB of RAM, running on Python 3.10.11 with PyTorch 2.0.0, Linux 5.15.0-71-generic, Ubuntu 22.04.2 LTS. For FATE framework, we are using federatedai/standalone_fate Docker image, running with Python 3.8.13 on Docker 23.0.2. Pivot is compiled from source using CMake 3.19.7, g++ 9.5.0, libboost 1.71.0, libscapi with git commit hash 1f70a88, and runs on a slurm cluster with AMD EPYC 7V13 64-Core Processor with the same number of cores as 2x AMD EPYC 7543 used for other algorithms. We conducted real distributed experiments on four distinct machines to evaluate the communication cost, the machines are equipped with AMD EPYC 7V13 64-Core Processors, 503Gi RAM, 1GbE NICs (Intel I350), and varying GPUs (AMD Instinct MI210 for $P_{1}$ and $P_{2}$, AMD Instinct MI100 for $P_{3}$ and $P_{4}$). The distributed experimental environment consisted of Python 3.9.12 and PyTorch 2.0.1+rocm5.4.2. We monitored network traffic from the network adapter before and after the experiments using ifconfig to measure the actual data size transferred between parties. We also recorded the network traffic size during idle server periods as background noise. Measurements were taken for 1 minute with a 1-minute gap between each measurement, totaling 5 times. The mean received data was 56.76 KB (std ±1.08 KB), and the mean sent data was 46.50 KB (std ±1.35 KB)
Distributed Implementations.
FedTree inherently supports distributed deployment through gRPC, a feature we utilize. In contrast, SplitNN, C-VFL, and GAL lack native distributed support. To address this, we implemented distributed support for these algorithms using torch.distributed. For the sparse matrix representation in C-VFL, we employ torch.sparse_coo_tensor.
@@
GAL Diao et al. [2022] MIT1
C-VFL Castiglia et al. [2022] MIT1
SecureBoost Cheng et al. [2021] Apache 2.02
Pivot Wu et al. [2020] CC BY-NC-ND 4.03
FedTree Li et al. [2023] Apache 2.02
@@ G.1 Communication Cost Details
In this subsection, we evaluate the total and maximum incoming/outgoing communication costs of various VFL algorithms over 50 epochs. This assessment is performed on four physically distributed machines, with results displayed in Figures 12, and 13. We observe real network adapter communication for both four and two-party settings with ifconfig. Additionally, theoretically estimated communication sizes are compared in Figure 14, and real-time performance in relation to increasing communication costs for four VFL algorithms is presented in Figure 15. The results yield following four observations.
Gradient-boosting algorithms, such as GAL and FedTree, tend to have small communication sizes when contrasted with NN-based methods like SplitNN and C-VFL. Yet, FedTree experiences marked communication overhead with high-dimensional datasets. C-VFL, due to its different pipeline from SplitNN, fails to reduce real communication cost, especially as it broadcasts compressed representations from all parties too all other parties. The large communication expenses in NN-based algorithms from the recurrent transmission of gradients and representations, contributing to SplitNN’s superior accuracy. The high communication costs of FedTree on high-dimensional datasets might be linked to the transmission of feature histograms and node information across parties.
@@
Generally, GAL is more efficient than FedTree, though FedTree sometimes has less incoming communication cost on low-dimensional datasets. The incoming communication cost includes residuals for GAL and histograms for FedTree. Residual size is dependent on the number of instances, while histogram size relates to the feature count. Hence, FedTree shows small incoming communication cost on low-dimensional datasets but more on high-dimensional datasets. However, FedTree’s outgoing communication cost is much larger than GAL, as it transmits both gradients and node information, resulting in its large overall communication overhead.
C-VFL struggles to reduce real communication costs via representation compression sent to the server, and it neither enhance the backward server-client communication. Initially, C-VFL’s incoming communication costs align closely with those of SplitNN. This is attributable to the method of estimating the compression ratio of representations by counting the ratio of non-zero elements. Though these representations appear sparse, they cannot be effectively transmitted without added overheads. Specifically, the coordinate format (COO) sparse matrix, as utilized in our experiments, demands even greater communication due to the need for additional indices. As a result, such compressions offer limited advantage in practical scenarios. Moreover, the consistent outgoing communication costs indicate that transmitting compressed aggregated models in C-VFL or broadcasting compressed representations does not display efficiency over transmitting uncompressed gradients of the cut layer. This revelation hints at avenues for potential refinement to minimize SplitNN’s backward communication expenses.
The real-time performance per communication round aligns with observations made for a fixed epoch. This is largely attributed to the dominance of communication per iteration over the number of iterations. GAL achieve similar accuracy levels with minimal communication overhead, a scenario where SplitNN and C-VFL might struggle to complete even a single iteration. However, when more communication is allowed, SplitNN typically delivers the most superior performance.
The estimated communication costs generally align with actual observations, with notable exceptions in C-VFL and FedTree when handling high-dimensional datasets. In the case of C-VFL, the estimation overlooks the overhead associated with transmitting sparse matrices. For FedTree on high-dimensional data, the reported communication cost likely arises from unaccounted transmission information related to nodes and histograms. Such discrepancies underscore the need for comprehensive testing on real distributed systems for VFL algorithms.
@@
Figure 12: Real total, max incoming and outgoing communication size across parties of VFL algorithms (50 global iterations)
Figure 13: Real total, max incoming and outgoing communication size across parties of VFL algorithms (two parties, 50 global iterations)
Figure 14: Estimated total communication size of VFL algorithms (50 global iterations)
Figure 15: Test performance as the increasing of communication cost (MiB)
@@ G.3 Training Time
Secondly, when comparing non-encryption methods, we find that split-based algorithms (SplitNN, FedTree) generally outperform ensemble-based algorithms (GAL) in terms of efficiency. This is primarily because split-based algorithms require each party to train a partial model, whereas ensemble-based algorithms mandate that each party train an entire model for ensemble purposes. This design characteristic also contributes to the lower communication costs associated with ensemble-based algorithms, as demonstrated in Figure 12.